# Estimation of Energy-dissipation Lower-bounds for Neuromorphic Learning-in-memory

### v3

@@ Results
We first create an abstraction of a memory model using the R-C circuit in Fig. 2(A) before generalizing it to other synaptic devices and bringing in concepts from stochastic thermodynamics [37, 38]. The capacitor $C$ acts as the memory element where the parameter stored in the memory at any instant of time $t$ is represented by the charge stored or the capacitor voltage $W(t)$. The parameter changes with time as stored charge discharges through a series of resistors $R_{0}$ and $R_{t}$ that model the interconnect and leakage resistance. Note that inductive effects are ignored in this derivation. If $R_{0}$ and $R_{t}$ are assumed to be constant, the voltage $W(t)$ decays from its initial value $W_{0}$ as
@@
$J(t)$ $=-C\frac{dW}{dt}=\frac{W_{0}}{(R_{0}+R_{t})}\exp\left[-\frac{t}{(R_{0}+R_{t})C}\right]$ (2)
$=\frac{W_{0}}{R_{0}}\exp\left[-\frac{t}{(R_{0}+R_{t})C}-\log\left(\frac{R_{0}+R_{t}}{R_{0}}\right)\right]$
as shown in Fig. 2(A)-(B). Note that if $R_{t}\rightarrow\infty$, then $J(t)\rightarrow 0$, as depicted in Fig. 2(B). However, if $R_{t}$ is arbitrarily modulated with respect to time, then obtaining a close-form solution of $J(t)$ is not always feasible. To make the analysis tractable, we discretize the modulation steps in intervals of $\Delta t$, and abstract the discrete time current $J_{n}^{0}=J(n\Delta t)$ as
@@
where $J_{max}=W_{0}/R_{0}$ is the maximum current that can flow out of the capacitor (determined by the interconnect resistance) and $E_{n}^{0}$ represents a time-varying energy-barrier that models both the charge decay and the modulation of the resistance $R_{t}$ in Fig. 2(A). This is depicted in Fig. 2(C) where the energy-barrier height $E_{n}^{0}$ keeps increasing with time (as $R_{t}\rightarrow\infty$) and accordingly $J_{n}^{0}$ decreases with time. In the quantization limit where the flow of current is due to discrete charged particles, $J_{n}^{0}/J_{max}$ can be viewed as a rate of transition (or escape) probability, a notion commonly used in Markovian dynamics [37, 39] and stochastic thermodynamics [40, 41, 42]. Equation 3 also uses coarse-graining procedure (like adiabatic elimination) from the stochastic thermodynamics literature [37] to ignore the fast-dynamics at the top of the energy-barrier and due to the reservoir (the sink). This description of physical memory is general enough and is applicable not to memory devices based on electrons stored on a capacitor, but to changes in conformations in molecular memory [43], DNA hybridization states [44, 45], or micro-filaments acting as memory states in a memristor. Eq. 3 can also be used to estimate the height of the energy barrier required to achieve a specific reliability in storage in steady-state, as specified by a precision variable $\delta=J_{\infty}^{0}/J_{max}$, and is given by $E_{\infty}^{0}=kT\log(1/\delta)$ [46]. In traditional AI systems using a conventional memory, this energy barrier is generally chosen to be high enough to prevent parameter leakage due to thermal fluctuations so that the parameter is retained in the memory throughout the entirety of training. For instance, in resistive random-access-memory (RRAM) devices, where the non-volatile state of the conductive filament between two electrodes determines the stored analog value [47], the energy barrier height can be as high as $1\text{\,}\mathrm{pJ}$ [48]. In charge-based devices like floating-gates or FeRAM, where the state of polarization determines the stored analog value [9, 49, 50], the energy barrier is typically around $10\text{\,}\mathrm{fJ}$ [50]. Therefore, an learning algorithm that adapts the stored weights in quantized steps $(\ldots,W_{n-1},W_{n},W_{n+1},\dots)$ so as to minimize some system-level loss-function consumes energy to overcome the energy-barrier $E_{\infty}^{0}$ for each of the parameter/memory updates. In LIM the energy barrier height $E^{0}_{n}$ that retains the learned parameter can be adapted, as shown in Fig. 2(C) and it was shown [35] that by matching the memory retention rates to the process of weight decay used in ML training energy efficiency could be significantly improved.
The memory model in Eq. 3 can be generalized to support differential updates (incremental and decremental) using an equivalent circuit model shown in Fig. 2(D). Here the parameters are stored as the difference between the voltages $W^{+}-W^{-}$ on the capacitors and the equivalent discharge currents are $J^{+}_{n}$ and $J^{-}_{n}$. Note that the net current flowing through the resistor $R_{n}$ is $J^{+}_{n}-J^{-}_{n}$. The capacitors are coupled to voltage sources $H_{1}$ and $H_{2}$, acting as equivalent chemical reservoirs and can shift the voltage levels $W^{+}$ and $W^{-}$ with respect to each other. When the system is in equilibrium
$J^{+}_{n}$ $=J^{-}_{n}=J^{0}_{n}.$ (4)
The circuit analogy can be generalized to other electronic devices using a band-diagram representation shown in Fig. 2(E). Note that while the currents $J^{+}_{n}$ and $J^{-}_{n}$ can flow between the reservoirs, energy is only dissipated through the resistors and is a function of the difference between the currents. We assume that $H_{1}$ and $H_{2}$ are ideal voltage sources (equivalent to a thermal reservoir with an infinite heat capacity) so the potential remains strictly constant during energy exchange. Energy transfer between the reservoirs and the capacitors is fully reversible, producing no entropy in the bath. Also note that as $R_{n}\rightarrow\infty$, $J^{+}_{n},J^{-}_{n}\rightarrow 0$; hence similar to Fig. 2(A) and (B), $W^{+}-W^{-}$ could be non-zero and encodes the stored parameter.
When the circuit is driven away from the equilibrium condition by changing the energy-levels as shown in Fig. 2(F), the respective currents change by
$J_{n}^{+}$ $=J_{n}^{0}+\frac{J_{n}}{2}$ (5)
$J_{n}^{-}$ $=J_{n}^{0}-\frac{J_{n}}{2},$
@@
where $J_{n}$ denotes the net current flowing across the barrier. If $E^{+}_{n}$ and $E^{-}_{n}$ denotes the effective barrier height for the respective node of the circuit, of which the energy differences $\Delta E_{n}=|E^{+}_{n}-E^{-}_{n}|$ encodes the stored memory. Then, assuming that the chemical reservoir is always at equilibrium [37, 39] at time-scales of the memory updates, $J_{n}$ (which represents probability flow) can be written as
$J_{n}$ $=J_{max}\cdot\left[\exp\left(-\frac{E^{+}_{n}}{kT}\right)-\exp\left(-\frac{E^{-}_{n}}{kT}.\right)\right]$ (6)
@@
$J_{n}=2J_{\text{max}}\exp\left(\frac{-E_{n}^{o}}{kT}\right)\left[\frac{\exp\left(\frac{-\Delta E_{n}}{kT}\right)-1}{\exp\left(\frac{-\Delta E_{n}}{kT}\right)+1}\right]$ (7)
and
$E_{n}^{0}$ $=kT\log\left[\frac{2J_{\text{max}}}{J_{n}}\cdot\frac{\exp\left(-\frac{\Delta E_{n}}{kT}\right)-1}{\exp\left(-\frac{\Delta E_{n}}{kT}\right)+1}\right].$ (8)
This relationship between $J_{n}$, $\Delta E_{n}$ and $E^{0}_{n}$ provides insight into how the barrier height ($E_{n}^{0}$) should change, or in turn, how the resistor $R_{n}$ should be modulated for a given memory update rate $J_{n}$ and external energy $\Delta E_{n}$.
In Fig. 3, we plot Eq. 8 for different values of the normalized update rate $r_{n}=\frac{J_{n}}{J_{\text{max}}}$. Along the $x$-axis the energy barrier is absent (i.e. for $E_{n}^{0}=0$) and the memory updates are driven directly by the presence of $\Delta E_{n}$. Along the $y$-axis the stored parameter reaches the trivial solution (i.e. $\Delta E_{n}=0$), hence the memory update rate is zero. Note that in Eq. 8, for $r_{n}=1$ and $\Delta E_{n}\rightarrow\infty$, when the update proceeds at maximum rate and the energy difference is initialized to infinitely large, we recover the absolute minimum energy barrier height equivalent to the celebrated Landauer’s limit $E^{0}_{n}=kT\log(2)$ [51]. Fig. 3A also shows the infeasible/inadmissible region where $r_{n}=\frac{\Delta J_{n}}{J_{\text{max}}}>1$, i.e. where $\Delta J_{n}$ exceeds the maximum update-rate $J_{\text{max}}$. During the process of learning, $\{(E_{n}^{0},\Delta E_{n})\}_{n=1}^{\infty}$ determines the computation rate $\{J_{n}\}_{n=1}^{\infty}$ according to the Eq. 8. Asymptotically, as $n\rightarrow\infty$, the learned parameters need to be retained, therefore memory leakage must be reduced by increasing the height of the energy barrier $E_{n}^{0}$, or $J_{n}$ must go to zero. To achieve this, different LIM algorithms will follow different trajectories in the $E^{0}_{n}$-vs-$\Delta E_{n}$ plot, as shown by $S_{1},S_{2},\ldots$ and $T_{1},T_{2},\ldots$ in Fig. 3(B). Asymptotically, trajectories $S$ and $T$ converge to $\Delta E_{S}$ and $\Delta E_{T}$, which represents the learned parameters (similar to the capacitor voltage in Fig. 2(B)).
@@
Fig. 4(A) demonstrates the dynamics of an LIM system with two parameters ($W_{x}$ and $W_{y}$) where learning involves minimizing an objective/loss function $L(\mathbf{W})$. Like the R-C example in Fig. 2(B), if the energy-barrier $E_{n}^{0}$ is held constant, the memory storing the parameters would decay from the initial state ${\bf W}^{0}$ to the trivial solution $\mathbf{W}^{trivial}$ following a fixed system energy contour $E(\mathbf{W})$ equivalent to $L(\mathbf{W})$. As shown in Fig. 4(A), at a finite time $T_{N}$ the energy-barrier modulates the update rate so that the effective energy contour $E(\mathbf{W},T_{N})$ is flat near the final solution. As a result, the memory updates would stop at a different state $\bf W$$^{*}$ and avoid decaying to the trivial solution $\mathbf{W}^{trivial}$. Note that the process of memory update (or $J_{n}$) is driven by the process of entropy production where ${\bf W}^{0}$ represents the initial low-entropy state. Hence in LIM, the process of learning (or memory updates) could follow different possible trajectories $T_{LIM}$ to the desired solution $\bf W$$^{*}$. The energy required to guide the process is used for modulating the respective energy barriers $E_{n}^{0}$ and the energy $\Delta E_{n}$ needed to update the stored parameters, according to the Eq. 8. Also for practical reasons, the final asymptotic (or retained) state needs to be within the neighborhood of $\bf W$$^{*}$ (to some predetermined precision). In the next section, we show the specific LIM trajectory could be chosen according to optimal memory consolidation dynamics.
@@
where $\beta>0$ and $kT$ are the conversion factors that has been introduced to convert algorithmic factors into physical factors. To simplify our analysis, we assume that the variance on the initial weights $\sigma^{2}(W_{d,0})$ can be ignored (due to deterministic initialization). We will use the assumptions described in the memory consolidation literature [29, 54], where $g_{d,n}$ are assumed to be uncorrelated with respect to each other in time and its variance is constant $\sigma_{g}^{2}$ (special case of signed update). Then,
@@
Since $E_{n}$ needs to be bounded (to model finite memory capacity), we follow the rationale provided in [29, 54] for an optimal memory consolidation schedule where $\epsilon_{k}$ decay is given by $\epsilon_{k}\sim(c/\sqrt{k})$. Substituting the optimal learning rate dynamics into Eq. 12 and using the fact that $0<\lambda c\ll 1$ [55, 56] we get
$\Delta E_{n}$ $=\beta kT\sigma_{g}^{2}c^{2}\sum_{k=1}^{n}\exp{\left(2\lambda c(\sqrt{n}-\sqrt{k})\right)}\frac{1}{k}$ (13)
$\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\sum_{k=1}^{n}\exp{\left(2\lambda c(\sqrt{n}-\sqrt{k})\right)}.$ (14)
Since the series in Eq. 14 are dominated by terms that are closer to $n$, let $\frac{p}{n}\ll 1$, $\quad p=n-k$, we can approximate $\sqrt{n}-\sqrt{k}$ using first order expansion as $\frac{p}{\sqrt{n}}$. Therefore, Eq. 14 can be rewritten as
$\Delta E_{n}$ $\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\sum_{p=0}^{n-1}\exp{\left(2\lambda c\frac{p}{\sqrt{n}}\right)}$ (15)
$\geq\beta kT\sigma_{g}^{2}\frac{c}{2\lambda}\frac{1}{\sqrt{n}}$ (16)
The detailed derivation can be found in Appendix section A.2, where we also show that the constants appeared in Eq. 14 are related to the precision of the memory retention $\delta$ according to $\beta\sigma_{g}^{2}\frac{c}{2\lambda}=1/\delta$. Note that Eq. 16 recovers the optimal learning rate dynamics $\epsilon_{n}$, which after combining with Eq. 8 leads to
$\Delta E_{n}\geq\left(\frac{kT}{\delta}\right)\frac{1}{\sqrt{n}}\quad$ (17)
and $E_{n}^{0}\geq kT\log\left[\frac{2J_{\text{max}}}{J_{n}}\cdot\frac{\exp\left(-1/\sqrt{n}\delta\right)-1}{\exp\left(-1/\sqrt{n}\delta\right)+1}\right]$ (18)
Eq.17-18 is model agnostic and connects the precision at which the parameters are updated and retained in memory during and at the end of training, and the resulting energy bounds. The update-rate $J_{n}$ models the speed of computation at a time-instant $n$, and $\delta$ models the precision of the memory update/retention, both of which are metrics associated with the update-wall. The learning-rate $\epsilon_{n}$ determines the dynamics of memory-consolidation [33, 57], and hence is a metric associated with the consolidation-wall. It is to be noted that, the precision modeled by $\delta$ is different from the $P$ bits of precision for the gradient of loss function. Here, $\delta$ denotes the precision of the memory retention/update at the final state. Thus, Eq. 17-18 is the key to derive and understand the energy-efficiency limits of LIM-based training.
@@
Eq. 17-18 shows that setting any two of the three parameters learning-rate $\epsilon_{n}$, update-rate $J_{n}$, energy barrier-height $\Delta E_{n}$, determines the third. This relationship can be used to estimate the total energy dissipation $E_{\text{Total}}$ for an LIM-based training. $E_{\text{Total}}$ is determined by the energy dissipated for building up the barrier $E_{\infty}^{0}$ at the end of training and the energy dissipated $\Delta E_{n}$ at each time instant $n$ to support the flow $J_{n}$.
@@
where $\Delta t$ denotes the duration of each discrete time step. Note that the expression for energy dissipation in equation 19 is aligned with the heat-flow expressions used in stochastic thermodynamics literature [37]. We also enforce constraints to ensure: (a) termination of training within a finite number of operations; and (b) the required memory retention at the end of training. These are specified as
$\textstyle M\sum_{n=1}^{\infty}J_{n}\Delta t=\#FLOPs$
 $\quad\mbox{ and }\quad E_{\infty}^{0}\geq kT\log(1/\delta)$ (20)
where #FLOPs\#FLOPs denotes the total computational workload need to complete a task. In Section .4, we use empirical trends based on reported metrics to estimate the #FLOPs\#FLOPs required. Note that unlike speed-limit theorems in stochastic thermodynamics [38], in Eq. 20 we are enforcing a task-completion within a finite time-horizon by choosing a specific schedule on the update rate $J_{n}$. We also constrain the update-rate schedule $r_{n}=\frac{J_{n}}{J_{\text{max}}}$ to be non-increasing, between zero and one, monotonic and asymptotically bounded (i.e. $r_{n}\overset{n\to\infty}{\longrightarrow}0$ and $\sum_{n=1}^{\infty}r_{n}\leq B,\quad B>0$). Note that these conditions complement the constraints imposed on update rate $r_{n}$ in Appendix section A.2, where we calibrate the hyper-parameters under the the adiabatic (slowest) bound on the update rate, while in practice, constraints in Eq. 20 enforce the total computation (#FLOPs\#FLOPs) remains bounded. Based on the constraints, a natural choice is a polynomial decay $r_{n}=n^{-(1+\gamma)}$ with $\gamma\geq 0$. This choice of $r_{n}$ ensures that $J_{n}\leq J_{\text{max}}$ and $\Delta E_{n}\geq 0$ for all $n$. Then Eq. 19 and Eq. 20 can be used to eliminate intermediate variables $J_{\text{max}}$ and $\Delta t$ to produce
@@
For LIM, the memory updates are driven by the energy stored as parameter in memory by Eq. 17 and shown in Fig. 3. We show in the Appendix subsectionA.3 that for $M$ parameters or memory elements, the total energy dissipation $E_{Total}$ for LIM under the optimal learning rate schedule $\epsilon_{n}\sim\frac{1}{\sqrt{n}}$ can be estimated as

### v4

@@ Results
We first create an abstraction of an analog memory model using an RC circuit in Fig. 2(A) before generalizing it to other synaptic devices. We emphasize that the RC circuit is not intended to capture the underlying physics, like metastability, of a typical memory device with well-defined energy barriers, but to illustrate two essential ingredients of any physical memory: (i) the presence of a stored state variable, and (ii) a dissipative path through which that state may be read out or modified. The variable resistance element provides an intuitive way to visualize how a dynamical energy barrier could modulate the rate of state evolution. Note that the discrete states in an analog memory arise because: (a) the stored values can only be differentiated from each other up to a certain precision, due to measurement artifacts; or (b) the inherent discreteness of the memory states, which could arise due to finite number of electrons in floating-gate memory [54, 61], or due to micro-filaments acting as memory states in a memristor [72] In the RC-model, the capacitor $C$ acts as the memory element where the parameter stored in the memory at any instant of time $t$ is represented by the charge stored or the capacitor voltage $W(t)$. The parameter changes with time as stored charge discharges through a series of resistors $R_{0}$ and $R_{t}$ that model the interconnect and leakage resistance. Note that inductive effects are ignored in this derivation. If $R_{0}$ and $R_{t}$ are assumed to be constant, the voltage $W(t)$ decays from its initial value $W_{0}$ as
@@
$aligned J(t)&=-C\frac{dW}{dt}=\frac{W_{0}}{(R_{0}+R_{t})}\exp\left[-\frac{t}{(R_{0}+R_{t})C}\right]\\ &=\frac{W_{0}}{R_{0}}\exp\left[-\frac{t}{(R_{0}+R_{t})C}-\log\left(\frac{R_{0}+R_{t}}{R_{0}}\right)\right]$ (2)
as shown in Fig. 2(A)-(B). Note that if $R_{t}\rightarrow\infty$, then $J(t)\rightarrow 0$, as depicted in Fig. 2(B). However, if $R_{t}$ is arbitrarily modulated with respect to time, then obtaining a close-form solution of $J(t)$ is not always feasible. To make the analysis tractable, we discretize the modulation steps in intervals of $\Delta t$, and abstract the discrete time current $J_{n}^{0}=J(n\Delta t)$ according to
@@
where $J_{max}$ is the maximum switching rate that is specific to different hardware substrates. A time-varying $E_{n}^{0}$ that modulates the update rate $J_{n}^{0}$ is functionally isomorphic to the variable resistance $R_{n}$ affecting the charge decay current in Fig. 2(A). This is depicted in Fig. 2(C) where the energy-barrier height $E_{n}^{0}$ keeps increasing with time (as $R_{n}\rightarrow\infty$) and accordingly $J_{n}^{0}$ decreases with time. $J_{n}^{0}/J_{max}$ can be viewed as a rate of transition (or escape) probability, a notion commonly used in Markovian dynamics [59, 67] and in stochastic thermodynamics [22, 74, 7]. Equation 3 also uses coarse-graining procedure (like adiabatic elimination) from the stochastic thermodynamics literature [59] to ignore the fast-dynamics at the top of the energy-barrier and due to the reservoir (the sink). Eq. 3 can also be used to estimate the height of the energy barrier required to achieve a specific reliability in storage in steady-state, as specified by a precision variable $\delta=J_{\infty}^{0}/J_{max}$, and is given by $E_{\infty}^{0}=kT\log(1/\delta)$ [4]. In traditional AI systems using a conventional memory, this energy barrier is generally chosen to be high enough to prevent parameter leakage due to thermal fluctuations so that the parameter is retained in the memory throughout the entirety of training. For instance, in RRAM devices, where the non-volatile state of the conductive filament between two electrodes determines the stored analog value [2], the energy barrier height can be as high as $1\text{\,}\mathrm{pJ}$ [77]. In charge-based devices like floating-gates or FeRAM, where the state of polarization determines the stored analog value [11, 55, Dünkel2017fefetCIMMVM], the energy barrier is typically around $10\text{\,}\mathrm{fJ}$ [Dünkel2017fefetCIMMVM]. Therefore, an learning algorithm that adapts the stored weights in quantized steps $(\ldots,W_{n-1},W_{n},W_{n+1},\dots)$ so as to minimize some system-level loss-function consumes energy to overcome the energy-barrier $E_{\infty}^{0}$ for each of the parameter/memory updates. In LIM the energy barrier height $E^{0}_{n}$ that retains the learned parameter can be adapted, as shown in Fig. 2(C) and it was shown [54] that by matching the memory retention rates to the process of weight decay used in ML training energy efficiency could be significantly improved.
The memory model in Eq. 3 can be generalized to support differential updates (incremental and decremental) using an equivalent circuit model shown in Fig. 2(D). Here the parameters are stored as the difference between the voltages $W^{+}-W^{-}$ on the capacitors and the equivalent discharge currents are $J^{+}_{n}$ and $J^{-}_{n}$. Note that the net current flowing through the resistor $R_{n}$ is $J^{+}_{n}-J^{-}_{n}$. The capacitors are coupled to voltage sources $H_{1}$ and $H_{2}$, acting as equivalent chemical reservoirs and can shift the voltage levels $W^{+}$ and $W^{-}$ with respect to each other. We assume that $H_{1}$ and $H_{2}$ are ideal voltage sources so the potential remains strictly constant during energy exchange. Energy transfer between the reservoirs and the capacitors is fully reversible, producing no entropy in the bath. Note that while the currents $J^{+}_{n}$ and $J^{-}_{n}$ can flow between the reservoirs, energy is only dissipated through the resistors and is a function of the difference between the currents. The bi-directional circuit analogy can similarly be mapped to the bi-potential band-diagram as shown Fig. 2(E), where the energetic difference across the energy barrier $E^{0}_{n}$ encodes the information stored. When no content is stored in the memory, the system is in equilibrium as {align} J^+_n = J^-_n=J^0_n. where the update rate in both directions are equal to each other, and thus no energy is dissipated. When memory is updated as shown in Fig. 2(E), an external field biases the energy landscape so that the top of the energy-levels on both sides no longer lie at equal energy. This energy asymmetry establishes a preferred state, which constitutes the stored information. Once the memory is updated, the effective finite barrier height $E^{+}_{n}$ and $E^{-}_{n}$ across the barrier $E^{0}_{n}$ determines the rate of thermally activated transitions and therefore the retention and update dynamics
$aligned J_{n}^{+}&=J_{n}^{0}+\frac{J_{n}}{2}\\ J_{n}^{-}&=J_{n}^{0}-\frac{J_{n}}{2},$ (4)
@@
where $J_{n}$ denotes the net update rate across the barrier. Also note that as $E^{0}_{n}\rightarrow\infty$, $J^{+}_{n},J^{-}_{n}\rightarrow 0$; hence similar to Fig. 2(A) and (B), the energetic difference across the barrier $\Delta E$ could be non-zero and encodes the stored parameter, which in turn causes the energy dissipation in the process of memory writes. Then, assuming that the chemical reservoir is always at equilibrium [59, 67] at time-scales of the memory updates, $J_{n}$ (which represents probability flow) can be written as
$aligned J_{n}&=J_{max}\cdot\left[\exp\left(-\frac{E^{+}_{n}}{kT}\right)-\exp\left(-\frac{E^{-}_{n}}{kT}.\right)\right]$ (5)
@@
$aligned J_{n}=2J_{\text{max}}\exp\left(\frac{-E_{n}^{o}}{kT}\right)\left[\frac{\exp\left(\frac{-\Delta E_{n}}{kT}\right)-1}{\exp\left(\frac{-\Delta E_{n}}{kT}\right)+1}\right]$ (6)
and {align} E_n^0 = kT log[2JmaxJn⋅exp(-ΔEnkT)-1exp(-ΔEnkT)+1]. This relationship between $J_{n}$, $\Delta E_{n}$ and $E^{0}_{n}$ provides insight into how the barrier height ($E_{n}^{0}$) should change, or in turn, how the resistor $R_{n}$ should be modulated for a given memory update rate $J_{n}$ and external energy $\Delta E_{n}$.
In Fig. 3, we plot Eq. 6 for different values of the normalized update rate $r_{n}=\frac{J_{n}}{J_{\text{max}}}$. Along the $x$-axis the energy barrier is absent (i.e. for $E_{n}^{0}=0$) and the memory updates are driven directly by the presence of $\Delta E_{n}$. Along the $y$-axis the stored parameter reaches the trivial solution (i.e. $\Delta E_{n}=0$), hence the memory update rate is zero. Note that in Eq. 6, in the limiting regime where the update rate approaches its maximal value ($r_{n}\!\to\!1$) and the energetic bias between states becomes very large ($\Delta E_{n}\!\to\!\infty$), the characteristic thermal energy scale associated with the minimum barrier height reduces to $E_{n}^{0}=kT\ln 2$. This scale reflects the fact that, in a thermally fluctuating environment, two memory states must be separated by a free-energy difference on the order of $kT$ in order to remain reliably distinguishable over the relevant timescale. The value $kT\ln 2$ coincides numerically with the Landauer bound for logically irreversible bit erasure [46]. It should be interpreted only as a consistency check on the underlying thermodynamic energy scale, rather than as a statement that the present analog retention process constitutes Landauer erasure. Fig. 3A also shows the infeasible/inadmissible region where $r_{n}=\frac{J_{n}}{J_{\text{max}}}>1$, i.e. where $J_{n}$ exceeds the maximum update-rate $J_{\text{max}}$. During the process of learning, $\{(E_{n}^{0},\Delta E_{n})\}_{n=1}^{\infty}$ determines the computation rate $\{J_{n}\}_{n=1}^{\infty}$ according to the Eq. 6. Asymptotically, as $n\rightarrow\infty$, the learned parameters need to be retained, therefore memory leakage must be reduced by increasing the height of the energy barrier $E_{n}^{0}$, or $J_{n}$ must go to zero. To achieve this, different LIM algorithms will follow different trajectories in the $E^{0}_{n}$-vs-$\Delta E_{n}$ plot, as shown by $S_{1},S_{2},\ldots$ and $T_{1},T_{2},\ldots$ in Fig. 3(B). Asymptotically, trajectories $S$ and $T$ converge to $\Delta E_{S}$ and $\Delta E_{T}$, which represents the learned parameters (similar to the capacitor voltage in Fig. 2(B)).
@@
Fig. 4(A) demonstrates the dynamics of an LIM system with two parameters ($W_{x}$ and $W_{y}$) where learning involves minimizing an objective/loss function $L(\mathbf{W})$. Like the R-C example in Fig. 2(B), if the energy-barrier $E_{n}^{0}$ is held constant, the memory storing the parameters would decay from the initial state ${\bf W}^{0}$ to the trivial solution $\mathbf{W}^{trivial}$ following a fixed system energy contour $E(\mathbf{W})$ equivalent to $L(\mathbf{W})$. As shown in Fig. 4(A), at a finite time $T_{n}$ the energy-barrier modulates the update rate so that the effective energy contour $E(\mathbf{W},T_{n})$ is flat near the final solution. As a result, the memory updates would stop at a different state $\bf W$$^{*}$ and avoid decaying to the trivial solution $\mathbf{W}^{trivial}$. Note that the process of memory update (or $J_{n}$) is driven by the process of entropy production where ${\bf W}^{0}$ represents the initial low-entropy state. Hence in LIM, the process of learning (or memory updates) could follow different possible trajectories $T_{LIM}$ to the desired solution $\bf W$$^{*}$. The energy required to guide the process is used for modulating the respective energy barriers $E_{n}^{0}$ and the energetic difference $\Delta E_{n}$ that encodes the stored parameters, according to the Eq. 6. Also for practical reasons, the final asymptotic (or retained) state needs to be within the neighborhood of $\bf W$$^{*}$ (to some predetermined precision). In the next section, we show the specific LIM trajectory could be chosen according to optimal memory consolidation dynamics.
@@
where $\beta>0$ and $kT$ are the conversion factors that has been introduced to convert algorithmic factors into physical factors. To simplify our analysis, we assume that the variance on the initial weights $\sigma^{2}(W_{d,0})$ can be ignored (due to deterministic initialization). We will use the assumptions described in the memory consolidation literature [23, 3], where $g_{d,n}$ are assumed to be uncorrelated with respect to each other in time and its variance is constant $\sigma_{g}^{2}$ (special case of signed update). It should be noted that this assumption is a common practice in theoretical machine learning [64] to understand the asymptotic convergence and for analytic tractability. However, in practice, many learning algorithms violate this assumption as described in the Discussion section.
Following Eq. 9
@@
Since $E_{n}$ needs to be bounded (to model finite memory capacity), we follow the rationale provided in [23, 3] for an optimal memory consolidation schedule where $\epsilon_{k}$ decay is given by $\epsilon_{k}\sim(c/\sqrt{k})$. Substituting the optimal learning rate dynamics into Eq. 10 and using the fact that $0<\lambda c\ll 1$ [30, 18] we get {align} ΔE_n = βkT σ_g^2 c^2∑_k=1^nexp(2λc(n - k))1k
βkT σ_g^2 c2n∑_k=1^nexp(2λc(n - k)). Since the series in Eq. 10 are dominated by terms that are closer to $n$, let $\frac{p}{n}\ll 1$, $\quad p=n-k$, we can approximate $\sqrt{n}-\sqrt{k}$ using first order expansion as $\frac{p}{\sqrt{n}}$. Therefore, Eq. 10 can be rewritten as {align} ΔE_n ≥βkT σ_g^2 c2n∑_p=0^n-1exp(2λcpn)
βkT σ_g^2 c2λ1n The detailed derivation can be found in Appendix section A.2, where we also show that the constants appeared in Eq. 10 are related to the precision of the memory $\delta$ according to $\beta\sigma_{g}^{2}\frac{c}{2\lambda}=1/\delta$. More specifically, $\delta$ denotes the precision at which a value can be stored in the memory or can be measured. The uncertainty in retention or measurement is reflected by the probability that noise (thermal fluctuations) causes an erroneous state transition or readout. Following the thermodynamic analysis of Brillouin [8] and Bennett [4], the minimal free energy required to reliably distinguish two physical states with error probability $\delta$ is equivalent to the minimal energy barrier height the memory must exhibit,
$E_{\infty}^{0}\geq kT\log(\frac{1}{\delta}).$ (11)
Also, based on Shannon information, resolving $P$ bits of information requires that the uncertainty in distinguishing the stored states be at most one part in $2^{P}$. Therefore, the probability of misidentifying a state must be no larger than $\delta\sim 2^{-P}$ and hence $\delta$ reflects the bit-precision $P$ of the model parameter.
Note that Eq. .2 recovers the optimal learning rate dynamics $\epsilon_{n}$, which after combining with Eq. 6 leads to {align} ΔE_n ≥(kTδ) 1n
and E_n^0 ≥kT log[2JmaxJn ⋅exp(-1/nδ)-1exp(-1/nδ)+1]
Eq..2-.2 is model agnostic and connects the precision at which the parameters are updated and retained in memory during and at the end of training, and the resulting energy bounds. The update-rate $J_{n}$ models the speed of computation at a time-instant $n$, and $\delta$ models the precision of the memory update/retention, both of which are metrics associated with the update-wall. The learning-rate $\epsilon_{n}$ determines the dynamics of memory-consolidation [52, 21], and hence is a metric associated with the consolidation-wall. It is to be noted that, the precision modeled by $\delta$ is different from the $P$ bits of precision for the gradient of loss function. Here, $\delta$ denotes the precision of the memory retention/update at the final state. Thus, Eq. .2-.2 is the key to derive and understand the energy-efficiency limits of LIM-based training.
@@
Eq. .2-.2 shows that setting any two of the three parameters learning-rate $\epsilon_{n}$, update-rate $J_{n}$, energy barrier-height $\Delta E_{n}$, determines the third. This relationship can be used to estimate the total energy dissipation $E_{\text{Total}}$ for an LIM-based training. $E_{\text{Total}}$ is determined by the energy dissipated for building up the barrier $E_{\infty}^{0}$ at the end of training and the energy dissipated $\Delta E_{n}$ at each time instant $n$ to support the flow $J_{n}$.
@@
where $\Delta t$ denotes the duration of each discrete time step. Note that the expression for energy dissipation in equation 12 is aligned with the heat-flow expressions used in stochastic thermodynamics literature [59]. We also enforce constraints to ensure: (a) termination of training within a finite number of operations; and (b) the required memory retention at the end of training. These are specified as {align} M ∑n=1 Jn Δt =#FLOPs
[…]
@@
For LIM, the memory updates are driven by the energy stored as parameter in memory by Eq. .2 and shown in Fig. 3. We show in the Appendix subsectionA.3 that for $M$ parameters or memory elements, the total energy dissipation $E_{Total}$ for LIM under the optimal learning rate schedule $\epsilon_{n}\sim\frac{1}{\sqrt{n}}$ can be estimated as